Update at 10:03:24(BST): The issue has now been resolved with this server.Phone support will be unavailable this evening from 5.30-8.00pm due to staffing issues. Please contact support@34sp.com for any support/billing queries.
Phones re-open at 8am, Thursday May 10th.
9th May 2012, at 5:35PM
10th May 2012, at 9:03AM
Update at 15:15:47(BST):
The network is now running normally. Apologies for any inconvenience caused.
Update at 12:28:02(BST):
This has now been resolved and is being monitored. Than you for your patience.
One of our clients servers has been compromised by hackers and is flooding our network. Engineers are liaising with the client at this time and we hope to mitigate the effects as soon as possible. Apologies for any inconvenience caused.
9th May 2012, at 12:09PM
9th May 2012, at 2:15PM
Update at 11:58:51(BST): The issue has now been resolved with this server.The server has been rebooted and our engineers and running some tests on the server there are no further issues.
We noticed an issue on the server and a reboot has been issued and a disk check has been started. Our engineers are currently resolving the issue and the server will be back online within the next few minutes.
We apologise for the inconvenience caused.
9th May 2012, at 11:44AM
9th May 2012, at 10:58AM
Update at 17:34:38(BST): Update at 20:51:39(BST): Services are now coming back up on the server.
Update at 20:47:53(BST): This server has now been rebooted
We are presently looking into an issue affecting this server.
7th May 2012, at 8:35PM
9th May 2012, at 4:34PM
Update at 16:31:56(BST): This issue has been resolved
7th May 2012, at 5:23PM
8th May 2012, at 3:31PM
Update at 16:32:32(BST): The changes seem to have worked and we are not seeing repeats of the ongoing issues on prohost8
4th May 2012, at 10:58AM
8th May 2012, at 3:32PM
Update at 19:03:25(BST): Update at 14:10:48(BST): We are still looking into the issue that is affecting the stability of this server.
3rd May 2012, at 10:04PM
15th May 2012, at 6:03PM
Update at 22:02:49(BST): The issue has now been resolved with this server.We are presently investigating a fault affecting webmail services on services that use our Atmail array.
3rd May 2012, at 9:31PM
3rd May 2012, at 9:02PM
Update at 12:29:18(BST):
This issue has been found and resolved. Apologies for any inconvenience caused and thank you for you patience during this issue.
Update at 09:31:54(BST): Update at 16:33:09(BST): This server is now behaving correctly.
2nd May 2012, at 11:11AM
9th May 2012, at 11:29AM
Update at 22:03:48(BST): Update at 14:04:21(BST): We now consider this case resolved.
2nd May 2012, at 11:10AM
3rd May 2012, at 9:03PM